OpenStudy (anonymous):

Problem #3: Find the total amount if you deposit $ 600 at a rate of 5% for two years compounded annually. Can you please just explain how to do this. I have another problem like this that I want to do on my own I just need to figure out how to do it first. Please help!

OpenStudy (anonymous):

you take the original value (600) times the percentage+1 (1.05) to the power of the time (two years) 600*1,05^2

OpenStudy (amistre64):

adjust r and n for appropriate periods \[B_0=B_0\] \[B_1=B_0(1+r)\] \[B_2=[B_0(1+r)](1+r)=B_0(1+r)^2\] \[B_3=[B_0(1+r)^2](1+r)=B_0(1+r)^3\] \[B_n=B_0(1+r)^n\]

OpenStudy (anonymous):

So then after one year it would be $631.13 and then after two years it would be $667?

OpenStudy (anonymous):

No, it would be 630 after 1 year and 661,5 after two

OpenStudy (amistre64):

600.00 630.00 661.50 694.58 etc ....

OpenStudy (anonymous):

How did you do it?

OpenStudy (anonymous):

I tried to do it the way that you had explained it.

OpenStudy (anonymous):

\[600\times(1,05)^2\]

OpenStudy (anonymous):

was that what you did?

OpenStudy (anonymous):

Ok So I am trying to do the next one can you check my answers to see if I am getting it right yet? Problem #4: Find the total amount if you deposit $800 at a rate of 8% for a total of four years compounded annually. Year1: 864 year2: 933.12 year 3: 989.10 year 4: 1,048.44

OpenStudy (anonymous):

You did the first one right but the rest is wrong, you shoul always take the original value (800) that doesn't change times 1+the percentage (1.08) that does not change either. The only thing that should be different between the years is what power you take the percentage to which is the amount of years (n): \[800\times(1.08)^n\]

OpenStudy (anonymous):

Year 2: 933.12 year3: 1,007.77 Year 4: 1,088.39 Does that look correct ?

OpenStudy (anonymous):

Yes!